Find the area between y=x^2 and x+y-2=0


1
Expert's answer
2021-12-20T08:07:14-0500
"y=x^2"

Substitute


"x+x^2-2=0"

"(x+2)(x-1)=0"

"x_1=-2, x_2=1"

"Point(-2, 4), \ Point(1,1)."


"x+y-2=0"

"y=-x+2"

"A=\displaystyle\int_{-2}^{1}(-x+2-x^2)dx"

"=\big[-\dfrac{x^3}{3}+2x-\dfrac{x^2}{2}\big]\begin{matrix}
 1 \\
 -2
\end{matrix}"

"=-\dfrac{1^3}{3}+2(1)-\dfrac{1^2}{2}-(\dfrac{(-2)^3}{3}+2(-2)-\dfrac{(-2)^2}{2})"

"=-\dfrac{1}{3}+2-\dfrac{1}{2}-\dfrac{8}{3}+4+2=\dfrac{9}{2}({units}^2)"



"Area=\dfrac{9}{2}" square units.